A motorboat traveling with the current went 72 miles in 3 hours. Against the current, the boat could only go 48 miles in the same amount of time. Find the rate of the boat in still water if the rate of the current is 4 miles per hour.
Answer by ewatrrr(24785)

r + c = 72/3 = 24
r - c = 48/3 = 18
2r = 42
r = 21 mph
IF
c = 4mph
r + 4 = 24
r = 20mph